# Sylvester domains and pro-$p$ groups

### v1

@@ 2.3. Skew power series rings
While we will not use them, one can define a left skew power series ring analogously. When $R$ is a Noetherian pseudocompact ring, our definition agrees with the one in [SV06] by seeing $S$ simultaneously as a right and left skew power series ring in the unique compatible way (see [SV06, Sec. 1]).
@@
In this paper, the ring $R$ will either be the completed group algebra $\mathbb{F}_{p}[\![N]\!]$ over $\mathbb{F}_{p}$ of a free pro-$p$ subgroup $N$ of a pro-$p$ group $G$ or a power series ring $\Lambda[\![t]\!]$ over a discrete $\mathbb{F}_{p}$-algebra $\Lambda$ for a fixed prime $p$. In both cases, $R$ is a complete ring whose topology is induced by a Hausdorff filtration $R_{k}$, where $R_{k}$ is either the intersection of the $k$-th power $I_{G}^{k}$ of the augmentation ideal of $\mathbb{F}_{p}[\![G]\!]$ with $\mathbb{F}_{p}[\![N]\!]$ or the principal ideal generated by $t^{k}$ of $\Lambda[\![t]\!]$. Following the nomenclature of [Laz65, Lab06], we have a filtration function
@@
which is submultiplicative, satisfies the ultrametric inequality and is such that $a_{i}\to 0$ if and only if $w(a_{i})\to 0$. In the $\mathbb{F}_{p}[\![N]\!]$ case, this filtration function is called a valuation in [EJZ13, Sec. 2.2], though we shall reserve the name “valuation” for a stronger class of filtration functions (see Section 4).
@@ 2.4. Universal division ring of fractions
Let $f:R\to S$ be a ring homomorphism. We say that $f$ is epic if for every ring $Q$ and homomorphisms $\alpha,\beta:S\to Q$, the equality $\alpha\circ f=\beta\circ f$ implies $\alpha=\beta$.
As in [Coh06, Chap. 7.2], we define an epic division $S$-ring as a division ring $\mathcal{D}$ together with an epic homomorphism $\varphi\colon S\to\mathcal{D}$. The condition on $\varphi$ to be epic is equivalent to the condition that the division closure of $\varphi(S)$ is equal to $\mathcal{D}$. Let $\varphi^{\prime}\colon S\to\mathcal{D}^{\prime}$ be another epic division $S$-ring. A subhomomorphism of epic division $S$-rings is a homomorphism $\psi\colon K\to\mathcal{D}^{\prime}$, where $K$ is a local subring of $\mathcal{D}$ containing $\varphi(S)$ with maximal ideal $\ker\psi$, such that $\psi\circ\varphi=\varphi^{\prime}$. Two subhomomorphisms are equivalent if their restriction to the intersection of their domains coincide and are again subhomomorphisms. An specialization $\mathcal{D}\to\mathcal{D}^{\prime}$ of epic division $S$-rings is an equivalence class of subhomomorphisms. The archetypal examples are $\mathcal{D}=\mathbb{Q}$ and $\mathcal{D}^{\prime}=\mathbb{F}_{p}$ for $S=\mathbb{Z}$, with the local subrings $K$ being the localization of $\mathbb{Z}$ at a prime $p$.
@@
The universal division $S$-ring $\mathcal{D}_{S}$ need not exist in general, even if $S$ is a domain ([Pas82]). One class of rings that possess a universal division ring of fractions is the class of Sylvester domains ([Coh06, Thm. 7.5.13]). To define a Sylvester domain, we consider Sylvester matrix rank functions and the inner rank. A Sylvester matrix rank function $\operatorname{rk}$ on an unital ring $S$ is a function defined on the set $\operatorname{Mat}(S)$ of all matrices over $S$ satisfying:
@@
    $\operatorname{rk}(AB)\leq\max\{\operatorname{rk}(A),\operatorname{rk}(B)\}$ for any pair of matrices $A$ and $B$ that can be multiplied;
@@
Since $\operatorname{rk}(\operatorname{Id_{n}})=n$ by (SMat3), one concludes that $\operatorname{rk}(A)\leq\min\{n,m\}$ for every matrix $A\in\operatorname{Mat}_{n\times m}(S)$ by property (SMat2) and the identities $A=\operatorname{Id}_{n}A=A\operatorname{Id}_{m}$. We include a proof of the following general fact for which we could find no reference:
@@
Every Sylvester matrix rank function defines a dimension function $\dim$ for finitely presented $S$-modules $M\simeq S^{m}/AS^{n}$ with $A\in\operatorname{Mat}_{n\times m}(S)$ through $\dim M=m-\operatorname{rk}(A)$. This dimension function is an example of a Sylvester module rank function (see [JZ20, Sec. 2.1]), and this correspondence gives a bijection between the sets of Sylvester matrix rank functions and Sylvester module rank functions.
@@
exists and defines a Sylvester matrix rank function on $S=\mathbb{F}_{p}[\![G]\!]$ that does not depend on the choice of the chain $U_{i}$ ([JZ19b, Prop. 11.2]). We denote the associated Sylvester module dimension function by $\dim_{G}$. Observe that for a finitely presented $\mathbb{F}_{p}[\![G]\!]$-module $M$ we have
@@
The inner rank $\operatorname{irk}_{S}(A)$ of a non-zero matrix $A\in\operatorname{Mat}_{n\times m}(S)$ is the smallest non-negative integer $k$ such that $A$ factors as a product $BC$ with $B\in\operatorname{Mat}_{n\times k}(S)$ and $C\in\operatorname{Mat}_{k\times m}(S)$. If the inner rank of a square $n\times n$ matrix $A$ is $n$, we call $A$ a full matrix. A ring $S$ is called a Sylvester domain if the inner rank $\operatorname{irk}$ is a Sylvester matrix rank function on $S$. Note that $\operatorname{irk}$ always satisfies the conditions (SMat1) and (SMat2), and every Sylvester matrix rank function $\operatorname{rk}$ on $S$ satisfies $\operatorname{rk}(A)\leq\operatorname{irk}(_{S}A)$ by (SMat2). Every Sylvester domain $S$ possesses a universal division ring of fractions $\mathcal{D}_{S}$ which satisfies another universal property:
Proposition 2.7 (cf. [Coh06, Thm. 7.5.13(e)]).
@@
If $\psi\colon S\to R$ is a homomorphism of rings and $\operatorname{rk}$ is a Sylvester matrix rank function on $R$, then the precomposition with $\psi$ defines a Sylvester matrix rank function on $S$, denoted $\psi^{\#}\operatorname{rk}$. Given a Sylvester matrix rank function $\operatorname{rk}^{\prime}$ on $S$, we say that $R$ is an envelope for $\operatorname{rk}$ if there exists $\psi$ and $\operatorname{rk}$ as above such that $\operatorname{rk}^{\prime}=\psi^{\#}\operatorname{rk}$.
We recall that a ring $\mathcal{U}$ is von Neumann regular if for every $a\in\mathcal{U}$ there exists $b\in\mathcal{U}$ such that $aba=a$. In particular, every division ring is von Neumann regular. In [JZ20], the first author proved the following:
Theorem 2.8 ([JZ20, Cor. 2.5]).
Let $S$ be a ring and $\operatorname{rk}$ be a Sylvester matrix rank function on $S$ with an envelope $\phi\colon R\to\mathcal{U}$ which is a von Neumann regular ring. Assume that:
@@
    For any finitely generated left or right $S$-submodule $M$ of $\mathcal{D}^{r}$ and any exact sequence $0\to I\to S^{n}\to M\to 0$, $I$ is a direct union of submodules isomorphic to $S^{k}$ where
@@ 3. Sylvester domains and completed group algebras
Proof.
We prove the lemma for left submodules, as the proof for right submodules is analogous. The main step of our proof is the following claim.
Claim 3.2.
@@
We provide two proofs of the claim. The first one is much easier but requires an additional condition that $\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}=\operatorname{rk}_{G}$. We notice that this is the case needed for the proof of Theorem 1.1. The second proof of Claim 3.2 does not require any additional hypothesis but uses a non-trivial result, proved in the appendix, that $\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ admits a Hausdorff ring topology such that the embedding $\mathbb{F}_{p}[\![G]\!]\to\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ is continuous.
First proof of Claim 3.2.
@@
be a chain of normal open subgroups of $G$ with trivial intersection. By [JZ24, Proposition 2.15], for every finitely generated left $\mathbb{F}_{p}[\![G]\!]$-module $N$,
@@
Second proof of Claim 3.2.
We can identify $\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ with the universal localization of $\mathbb{F}_{p}[\![G]\!]$ at the set of all full matrices. Thus, by Theorem A.1, $\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ admits a Hausdorff ring topology such that the embedding $\mathbb{F}_{p}[\![G]\!]\to\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ is continuous. In particular, $I$ is closed in $\mathbb{F}_{p}[\![G]\!]^{n}$. ∎
@@
Let $J=I/I_{G}I$, where $I_{G}$ is the augmentation ideal of $\mathbb{F}_{p}[\![G]\!]$. Since $I$ is the direct union of submodules isomorphic to $\mathbb{F}_{p}[\![G]\!]^{k}$, the $\mathbb{F}_{p}$-vector space $J$ must be isomorphic to the direct union of subspaces of dimension at most $k$. Therefore $\dim_{\mathbb{F}_{p}}J\leq k$ and there exists $k$ such that $J=I_{k}+I_{G}I$. Since $I$ is closed, it is finitely generated. Hence, by Nakayama’s lemma, $I$ must be equal to $I_{k}$. ∎
The following observation is an immediate consequence of the previous lemma:
@@
Apply Lemma 3.1 with $M$ being the augmentation ideal of $\mathbb{F}_{p}[\![G]\!]$. ∎
@@
(2) $\begin{gathered}\tau(g_{1}g_{2},g_{3})\sigma(g_{3})(\tau(g_{1},g_{2}))=\tau(g_{1},g_{2}g_{3})\tau(g_{2},g_{3})\,,\\ \sigma(g_{2})(\sigma(g_{1})(r))=\tau(g_{1},g_{2})^{-1}\sigma(g_{1}g_{2})(r)\tau(g_{1},g_{2})\,,\end{gathered}$
for every $g_{1},g_{2},g_{3}\in G$ and $r\in R$ – this follows directly from the associativity of the product in $S$ ([Pas13, Lem. 1.1]). Since $u_{g}$ are invertible in $S$, $\{u_{g}\mid g\in G\}$ is a free basis of $S$ as a left $R$-module.
@@
(c) Let $T$ be a transversal of $U$ in $G$. Then $T=\{g_{1},\ldots,g_{n}\}$ is a $\mathcal{Q}$-basis of $\mathcal{Q}*G/U$. for every $i=1,\ldots,n$ we define $\gamma_{i}\in E$ such that $\gamma_{i}(g_{i})=\delta_{ij}g_{j}$. We identify the elements of the ring $\mathcal{Q}*G/U$ with its images in $E$
Claim 3.5.
We have that $E=\bigoplus_{i=1}^{n}(\mathcal{Q}*G/U)\circ\gamma_{i}$.
Proof.
It is enough to show that if $\alpha_{1},\ldots,\alpha_{n}\in\mathcal{Q}*G/U$ and $\alpha=\sum_{i=1}^{n}\alpha_{i}\circ\gamma_{i}=0$, then $\alpha_{1}=\ldots=\alpha_{n}=0$. This follows from the equality $\alpha(g_{i})=\alpha_{i}g_{i}$. ∎
Claim 3.6.
We have that $E=\bigoplus_{i=1}^{n}\gamma_{i}\circ(\mathcal{Q}*G/U)$.
Proof.
It is enough to show that if $\alpha_{1},\ldots,\alpha_{n}\in\mathcal{Q}*G/U$ and $\alpha=\sum_{i=1}^{n}\gamma_{i}\circ\alpha_{i}=0$, then $\alpha_{1}=\ldots=\alpha_{n}=0$. Write $\alpha_{i}=\sum_{j=1}^{n}\alpha_{ij}g_{j}$. Then
@@
Now it is cleat that if all $\alpha(g_{k})=0$ then for every $i,j$, $\alpha_{ij}=0$. ∎
(d) Just observe that the map $\mathcal{Q}\otimes_{\mathbb{F}_{p}[\![U]\!]}\mathbb{F}_{p}[\![G]\!]\to\mathcal{Q}*G/U$ sending $a\otimes b$ to $ab$ is an isomorphism of right $\mathbb{F}_{p}[\![G]\!]$-modules.
@@
by Proposition 2.9. Since $\mathcal{Q}$ is the direct limit of it’s finitely generated submodules, we conclude that $\operatorname{Tor}_{1}^{\mathbb{F}_{p}[\![U]\!]}(\mathcal{Q},\mathcal{Q})=0$.
@@
to conclude that $I$ must be finitely presented and hence of type $FP_{\infty}$ over $\mathbb{F}_{p}[\![U]\!]$. Therefore, we conclude that
$\operatorname{Tor}_{1}^{\mathbb{F}_{p}[\![U]\!]}(I,\mathbb{F}_{p})\simeq\operatorname{Tor}_{1}^{\mathbb{F}_{p}[\![G]\!]}(I\otimes_{\mathbb{F}_{p}[\![U]\!]}\mathbb{F}_{p}[\![G]\!],\mathbb{F}_{p})=0\,.$
Since $\mathbb{F}_{p}[\![U]\!]$ is local and $I$ is finitely generated, $I$ must be projective over $\mathbb{F}_{p}[\![U]\!]$ and hence free (and of finite rank). By Theorem 2.8, $\mathbb{F}_{p}[\![U]\!]$ is a Sylvester domain and its division closure $\mathcal{Q}_{0}$ inside $\mathcal{Q}$ is $\mathbb{F}_{p}[\![U]\!]$-isomorphic to it’s universal division ring of fractions.
@@
By Proposition 3.3, $U$ is of cohomological dimension 2. By Serre’s theorem [Ser65] (see also [Har90]), $G$ is also of cohomological dimension 2. Thus, since $M$ is a finitely presented submodule of $(\mathcal{Q}*G/U)^{p}$, there is an exact sequence of $\operatorname{Tor}$ groups:

### v2

@@ 2.3. Skew power series rings
While we will not use it, one can define a left skew power series ring analogously. When $R$ is a Noetherian pseudocompact ring, our definition agrees with the one in [45] by seeing $S$ simultaneously as a right and left skew power series ring in the unique compatible way (see [45, Sec. 1]).
@@
In this paper, the ring $R$ will either be the completed group algebra $\mathbb{F}_{p}[\![N]\!]$ over $\mathbb{F}_{p}$ of a free pro-$p$ subgroup $N$ of a pro-$p$ group $G$ or a power series ring $\Lambda[\![t]\!]$ over a discrete $\mathbb{F}_{p}$-algebra $\Lambda$ for a fixed prime $p$. In both cases, $R$ is a complete ring whose topology is induced by a Hausdorff filtration $R_{k}$, where $R_{k}$ is either the intersection of the $k$-th power $I_{G}^{k}$ of the augmentation ideal of $\mathbb{F}_{p}[\![G]\!]$ with $\mathbb{F}_{p}[\![N]\!]$ or the principal ideal generated by $t^{k}$ of $\Lambda[\![t]\!]$. Following the nomenclature of [31, 28], we have a filtration function
@@
which is submultiplicative, satisfies the ultrametric inequality and is such that $a_{i}\to 0$ if and only if $w(a_{i})\to 0$. In the $\mathbb{F}_{p}[\![N]\!]$ case, this filtration function is called a valuation in [12, Sec. 2.2], though we shall reserve the name “valuation” for a stronger class of filtration functions (see Section 4).
@@ 2.4. Universal division ring of fractions
Let $f:R\to S$ be a ring homomorphism. We say that $f$ is epic if for every ring $Q$ and homomorphisms $\alpha,\beta:S\to Q$, the equality $\alpha\circ f=\beta\circ f$ implies $\alpha=\beta$ (i.e., the natural map $\operatorname{Hom}_{\text{Ring}}(S,Q)\to\operatorname{Hom}_{\text{Ring}}(R,Q)$ is injective for all rings $Q$).
As in [10, Chap. 7.2], we define an epic division $S$-ring as a division ring $\mathcal{D}$ together with an epic homomorphism $\varphi\colon S\to\mathcal{D}$. The condition on $\varphi$ to be epic is equivalent to the condition that the division closure of $\varphi(S)$ is equal to $\mathcal{D}$ ([10, Cor. 7.2.2]). Let $\varphi^{\prime}\colon S\to\mathcal{D}^{\prime}$ be another epic division $S$-ring. A subhomomorphism of epic division $S$-rings is a homomorphism $\psi\colon K\to\mathcal{D}^{\prime}$, where $K$ is a local subring of $\mathcal{D}$ containing $\varphi(S)$ with maximal ideal $\ker\psi$, such that $\psi\circ\varphi=\varphi^{\prime}$. Two subhomomorphisms $\psi_{1}\colon K_{1}\to\mathcal{D}$ and $\psi_{2}\colon K_{2}\to\mathcal{D}$ are equivalent if there is a subring $K_{0}$ of $\mathcal{D}$ contained in $K_{1}\cap K_{2}$ such that $\psi_{1}$ and $\psi_{2}$ agree on $K_{0}$ and it is local with maximal ideal $\ker(\psi_{1})\cap K_{0}=\ker(\psi_{2})\cap K_{0}$. A specialization $\mathcal{D}\to\mathcal{D}^{\prime}$ of epic division $S$-rings is an equivalence class of subhomomorphisms. The archetypal examples are $\mathcal{D}=\mathbb{Q}$ and $\mathcal{D}^{\prime}=\mathbb{F}_{p}$ for $S=\mathbb{Z}$, with the local subrings $K$ being the localization of $\mathbb{Z}$ at a prime $p$.
@@
The universal division $S$-ring $\mathcal{D}_{S}$ need not exist in general, even if $S$ is a domain ([37]). One class of rings that possess a universal division ring of fractions is the class of Sylvester domains ([10, Thm. 7.5.13]). To define a Sylvester domain, we consider Sylvester matrix rank functions and the inner rank. A Sylvester matrix rank function $\operatorname{rk}$ on an unital ring $S$ is a non-negative real-valued function defined on the set $\operatorname{Mat}(S)$ of all matrices over $S$ satisfying:
@@
    $\operatorname{rk}(AB)\leq\min\{\operatorname{rk}(A),\operatorname{rk}(B)\}$ for any pair of matrices $A$ and $B$ that can be multiplied;
@@
Since $\operatorname{rk}(\operatorname{Id_{n}})=n$ by (SMat3), one concludes that $\operatorname{rk}(A)\leq\min\{n,m\}$ for every matrix $A\in\operatorname{Mat}_{n\times m}(S)$ by property (SMat2) and the identities $A=\operatorname{Id}_{n}A=A\operatorname{Id}_{m}$. If $S$ is a division ring, then any matrix can be put in row-echelon or column-echelon form by multiplication with invertible matrices, an operation that doesn’t change the rank by (SMat2). It follows then from (SMat4) that there is a unique Sylvester matrix rank function on $S$, given by the number of linearly independent rows (with a left $S$-action) or columns (with a right $S$-action). We include a proof of the following general fact for which we could find no reference:
@@
Every Sylvester matrix rank function defines a dimension function $\dim$ for finitely presented $S$-modules $M\simeq S^{m}/AS^{n}$ with $A\in\operatorname{Mat}_{n\times m}(S)$ through $\dim M=m-\operatorname{rk}(A)$. This dimension function is an example of a Sylvester module rank function (see [24, Sec. 2.1]), and this correspondence gives a bijection between the sets of Sylvester matrix rank functions and Sylvester module rank functions.
@@
exists, where $A_{i}$ is seen as a linear operator $\mathbb{F}_{p}^{m|G\colon U_{i}|}\to\mathbb{F}_{p}^{n|G\colon U_{i}|}$ after some choice of basis for $\mathbb{F}_{p}[G/U_{i}]$, and defines a Sylvester matrix rank function on $S=\mathbb{F}_{p}[\![G]\!]$ that does not depend on the choice of the chain $U_{i}$ ([22, Prop. 11.2]). We denote the associated Sylvester module dimension function by $\dim_{G}$. Observe that for a finitely presented $\mathbb{F}_{p}[\![G]\!]$-module $M$ we have
@@
The inner rank $\operatorname{irk}_{S}(A)$ of a non-zero matrix $A\in\operatorname{Mat}_{n\times m}(S)$ is the smallest non-negative integer $k$ such that $A$ factors as a product $BC$ with $B\in\operatorname{Mat}_{n\times k}(S)$ and $C\in\operatorname{Mat}_{k\times m}(S)$. If the inner rank of a square $n\times n$ matrix $A$ is $n$, we call $A$ a full matrix. A ring $S$ is called a Sylvester domain if the inner rank $\operatorname{irk}$ is a Sylvester matrix rank function on $S$. Note that $\operatorname{irk}$ always satisfies the conditions (SMat1) and (SMat2), and every Sylvester matrix rank function $\operatorname{rk}$ on $S$ satisfies $\operatorname{rk}(A)\leq\operatorname{irk}(_{S}A)$ by (SMat2). In fact, $\operatorname{irk}$ satisfies an even stronger property than (SMat1): $\operatorname{irk}A=0$ if and only if $A$ is the zero matrix. Every Sylvester domain $S$ possesses a universal division ring of fractions $\mathcal{D}_{S}$ such that $\operatorname{irk}=\operatorname{rk}_{\mathcal{D}_{S}}$. For the construction of $\mathcal{D}_{S}$, see [10, Sec. 7.4 and 7.5]. Moreover, $\mathcal{D}_{S}$ satisfies another universal property:
Proposition 2.7 (cf. [10, Thm. 7.5.13(e)]).
@@
If $\psi\colon S\to R$ is a homomorphism of rings and $\operatorname{rk}$ is a Sylvester matrix rank function on $R$, then the precomposition with $\psi$ defines a Sylvester matrix rank function on $S$, denoted $\psi^{\#}\operatorname{rk}$. Given a Sylvester matrix rank function $\operatorname{rk}^{\prime}$ on $S$, we say that $R$ is an envelope for $\operatorname{rk}^{\prime}$ if there exists $\psi$ and $\operatorname{rk}$ as above such that $\operatorname{rk}^{\prime}=\psi^{\#}\operatorname{rk}$.
We recall that a ring $\mathcal{U}$ is von Neumann regular if for every $a\in\mathcal{U}$ there exists $b\in\mathcal{U}$ such that $aba=a$. In particular, every division ring is von Neumann regular. In [24], the first author proved the following:
Theorem 2.8 ([24, Cor. 2.5]).
Let $S$ be a ring and $\operatorname{rk}$ be a Sylvester matrix rank function on $S$ with an envelope $\phi\colon S\to\mathcal{U}$ which is a von Neumann regular ring. Assume that:
@@
    For any finitely generated left or right $S$-submodule $M$ of $\mathcal{D}^{r}$ and any exact sequence $0\to I\to S^{n}\to M\to 0$, $I$ is a (set-theoretic) union of submodules isomorphic to $S^{k}$ where
@@
Remark 2.10.
A big technical difficulty in proving Theorem 1.1 is establishing the validity of the condition (2) of Theorem 2.8 for the embedding $\mathbb{F}_{p}[\![G]\!]\to\mathcal{Q}$ we construct in Section 5. If $G$ is a non-trivial pro-$p$ $p$-adic analytic group and $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain, then $\operatorname{cd}G\leq 2$ by Proposition 3.3 and hence $G\simeq\mathbb{Z}_{p}\rtimes\mathbb{Z}_{p}$ or $G\simeq\mathbb{Z}_{p}$. In particular, $G$ is free-by-cyclic and Theorem 1.1 applies, showing that those are precisely the $p$-adic Lie groups $G$ for which $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain. However, for any torsion-free $p$-adic Lie group $G$ the completed group algebra $\mathbb{F}_{p}[\![G]\!]$ is a Noetherian domain, and hence it possesses a classical Ore ring of fractions $\mathcal{Q}$. For $G=\mathbb{Z}_{p}\rtimes\mathbb{Z}_{p}$ or $G=\mathbb{Z}_{p}$ it is straightforward to see that the map $\mathbb{F}_{p}[\![G]\!]\to\mathcal{Q}$ satisfies the conditions of Theorem 2.8. Since $\mathcal{Q}$ is flat over $\mathbb{F}_{p}[\![G]\!]$, we have $\operatorname{Tor}_{1}^{\mathbb{F}_{p}[\![G]\!]}(\mathcal{Q},\mathcal{Q})=0$. Moreover, any finitely generated $G$-submodule $M$ of $\mathcal{Q}$ is isomorphic to a finitely generated $G$-submodule of $\mathbb{F}_{p}[\![G]\!]$ by the Ore condition. Observe that any submodule of a finitely generated $\mathbb{F}_{p}[\![G]\!]$ is finitely generated and hence closed and profinite. Since $\mathbb{F}_{p}[\![G]\!]$ has global dimension at most $2$, by [8, Rem. after Thm. 3.5] we have
$\mathfrak{Tor}_{1}^{\mathbb{F}_{p}[\![G]\!]}(I,\mathbb{F}_{p})\simeq\operatorname{Tor}_{1}^{\mathbb{F}_{p}[\![G]\!]}(I,\mathbb{F}_{p})\simeq\operatorname{Tor}_{2}^{\mathbb{F}_{p}[\![G]\!]}(M,\mathbb{F}_{p})\leq\operatorname{Tor}_{2}^{\mathbb{F}_{p}[\![G]\!]}(\mathbb{F}_{p}[\![G]\!],\mathbb{F}_{p})=0\,$,
where $\mathfrak{Tor}$ denotes the derived functor of the completed tensor product. By [8, Prop. 3.1], $I$ is projective, and thus $I$ is a free $\mathbb{F}_{p}[\![G]\!]$-module since the latter is a local ring. For the fact that $\mathbb{F}_{p}[\![G]\!]$ is a local Noetherian domain of global dimension $\operatorname{cd}G$, see [4]. The fact that $\operatorname{rk}_{G}=\operatorname{rk}_{\mathcal{Q}}$ is also true for any $p$-adic analytic group, and is a result of M. Harris ([16, Lem. 1.10.1]).
@@ 3. Sylvester domains and completed group algebras
The main step for obtaining this result is:
Lemma 3.2.
@@
We provide two proofs of the lemma. The first one is much easier but requires an additional condition that $\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}=\operatorname{rk}_{G}$. We notice that this is the case needed for the proof of Theorem 1.1. The second proof of Lemma 3.2 does not require any additional hypothesis but uses a non-trivial result, proved in the appendix, that $\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ admits a Hausdorff ring topology such that the embedding $\mathbb{F}_{p}[\![G]\!]\to\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ is continuous.
First proof of Lemma 3.2.
@@
be a chain of normal open subgroups of $G$ with trivial intersection. By [26, Proposition 2.15], for every finitely generated left $\mathbb{F}_{p}[\![G]\!]$-module $N$,
@@
Second proof of Lemma 3.2.
We can identify $\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ with the universal localization of $\mathbb{F}_{p}[\![G]\!]$ at the set of all full matrices. Thus, by Theorem A.1, $\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ admits a Hausdorff ring topology such that the embedding $\mathbb{F}_{p}[\![G]\!]\to\mathcal{D}_{\mathbb{F}_{p}[\![G]\!]}$ is continuous. Hence, if $m_{1}\,,\ldots\,,m_{n}$ is a set of generators for $M$, the map $\mathbb{F}_{p}[\![G]\!]^{n}\to M$ sending $(d_{1},\ldots,d_{n})$ to $d_{1}m_{1}+\cdots+d_{n}m_{n}$ is also continuous. In particular, the kernel $I$ of this map is closed in $\mathbb{F}_{p}[\![G]\!]^{n}$. ∎
Proof of Lemma 3.1.
@@
Let $J=I/I_{G}I$, where $I_{G}$ is the augmentation ideal of $\mathbb{F}_{p}[\![G]\!]$. Since $I$ is the direct union of submodules isomorphic to $\mathbb{F}_{p}[\![G]\!]^{k}$, the $\mathbb{F}_{p}$-vector space $J$ must be isomorphic to the direct union of subspaces of dimension at most $k$. Therefore $\dim_{\mathbb{F}_{p}}J\leq k$ and there exists $i$ such that $J=I_{i}+I_{G}I$. Since $I$ is closed by Lemma 3.2, it is finitely generated. Hence, by Nakayama’s lemma, $I$ must be equal to $I_{i}$. ∎
The following observation is an immediate consequence of the previous lemmas:
@@
Apply Lemma 3.1 with $M$ being the augmentation ideal of $\mathbb{F}_{p}[\![G]\!]$. Observe that $G$ is finitely presented if and only if
$H_{2}(G,\mathbb{F}_{p})=\operatorname{Tor}_{2}^{\mathbb{F}_{p}[\![G]\!]}(\mathbb{F}_{p},\mathbb{F}_{p})\simeq\operatorname{Tor}_{1}^{\mathbb{F}_{p}[\![G]\!]}(M,\mathbb{F}_{p})\leq\operatorname{Tor}_{0}^{\mathbb{F}_{p}[\![G]\!]}(I,\mathbb{F}_{p})$
is finite, that is, if $I$ is finitely generated. ∎
@@
[…]
for every $g_{1},g_{2},g_{3}\in G$ and $r\in R$ – this follows directly from the associativity of the product in $S$ ([38, Lem. 1.1]). Since $u_{g}$ are invertible in $S$, $\{u_{g}\mid g\in G\}$ is a free basis of $S$ as a left $R$-module.
@@
(c) Let $T$ be a transversal of $U$ in $G$. Then $T=\{g_{1},\ldots,g_{n}\}$ is a $\mathcal{Q}$-basis of $\mathcal{Q}*G/U$. for every $i=1,\ldots,n$ we define $\gamma_{i}\in E$ such that $\gamma_{i}(g_{i})=\delta_{ij}g_{j}$. We identify the elements of the ring $\mathcal{Q}*G/U$ with its images in $E$.
We have that $E=\bigoplus_{i=1}^{n}(\mathcal{Q}*G/U)\circ\gamma_{i}$. To see it, it is enough to show that if $\alpha_{1},\ldots,\alpha_{n}\in\mathcal{Q}*G/U$ and $\alpha=\sum_{i=1}^{n}\alpha_{i}\circ\gamma_{i}=0$, then $\alpha_{1}=\ldots=\alpha_{n}=0$. This follows from the equality $\alpha(g_{i})=\alpha_{i}g_{i}$. We also have that $E=\bigoplus_{i=1}^{n}\gamma_{i}\circ(\mathcal{Q}*G/U)$, for which it is enough to show that if $\alpha_{1},\ldots,\alpha_{n}\in\mathcal{Q}*G/U$ and $\alpha=\sum_{i=1}^{n}\gamma_{i}\circ\alpha_{i}=0$, then $\alpha_{1}=\ldots=\alpha_{n}=0$. Write $\alpha_{i}=\sum_{j=1}^{n}\alpha_{ij}g_{j}$. Then
@@
Now it is clear that if all $\alpha(g_{k})=0$ then for every $i,j$, $\alpha_{ij}=0$.
(d) Just observe that the map $\mathcal{Q}\otimes_{\mathbb{F}_{p}[\![U]\!]}\mathbb{F}_{p}[\![G]\!]\to\mathcal{Q}*G/U$ sending $a\otimes b$ to $ab$ is an isomorphism of right $\mathbb{F}_{p}[\![G]\!]$-modules, whose inverse is given by sending $q_{1}g_{1}+\cdots+q_{n}g_{n}$ to $q_{1}\otimes g_{1}+\cdots+q_{n}\otimes g_{n}$.
@@
by Proposition 2.9. Since $\mathcal{Q}$ is the direct limit of its finitely generated submodules, we conclude that $\operatorname{Tor}_{1}^{\mathbb{F}_{p}[\![U]\!]}(\mathcal{Q},\mathcal{Q})=0$.
@@
to conclude that $I$ must be finitely presented and hence of type $FP_{\infty}$ over $\mathbb{F}_{p}[\![U]\!]$. Therefore, by [8, Lem. 2.1] we conclude that
$\mathfrak{Tor}_{1}^{\mathbb{F}_{p}[\![U]\!]}(I,\mathbb{F}_{p})\simeq\operatorname{Tor}_{1}^{\mathbb{F}_{p}[\![U]\!]}(I,\mathbb{F}_{p})\simeq\operatorname{Tor}_{1}^{\mathbb{F}_{p}[\![G]\!]}(I\otimes_{\mathbb{F}_{p}[\![U]\!]}\mathbb{F}_{p}[\![G]\!],\mathbb{F}_{p})=0\,$,
where $\mathfrak{Tor}$ is the derived functor of the completed tensor product of profinite modules. Since $I$ is profinite, $I$ must be projective over $\mathbb{F}_{p}[\![U]\!]$ by [8, Prop. 3.1] and hence free (and of finite rank) by Kaplansky’s theorem on projective modules over local rings. By Theorem 2.8, $\mathbb{F}_{p}[\![U]\!]$ is a Sylvester domain and its division closure $\mathcal{Q}_{0}$ inside $\mathcal{Q}$ is $\mathbb{F}_{p}[\![U]\!]$-isomorphic to it’s universal division ring of fractions.
@@
By Proposition 3.3, $U$ is of cohomological dimension 2. By Serre’s theorem [47] (see also [15]), $G$ is also of cohomological dimension 2. Thus, since $M$ is a finitely presented submodule of $(\mathcal{Q}*G/U)^{p}$, there is an exact sequence of $\operatorname{Tor}$ groups: